# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v1

@@ 7 Conclusion
Appendix A Notation
We denote $T$ as the total number of rounds and $t\in[T]$ as the current round. We denote $N$ as the total number of items.
@@ Appendix A Notation
We offer a Table A.1 for convenient reference.
Table A.1: Symbols
@@
$\tilde{R}_{t}(S)$ $:=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{ti})}$, optimistic expected revenue for the assortment $S$ at round $t$
Appendix B Properties of MNL function
@@
B.1 Utility for Outside Option: $v_{0}=\Theta(1)$ is Common in Contextual MNL Bandits
In this subsection, we explain why the assumption that $v_{0}=\Theta(1)$ is made without loss of generality. Let the original feature vectors be $x^{\prime}_{ti}\in\mathbb{R}^{d}$ for every item $i\in[N]$. Suppose that a context for the outside option $x_{t0}^{\prime}$ is given and the he probability of choosing any item $i\in S_{t}\cup\{0\}$ is defined as
$p_{t}(i|S_{t},\mathbf{w}^{\star})=\frac{\exp((x^{\prime}_{ti})^{\top}\mathbf{w}^{\star})}{\sum_{j\in S_{t}\cup\{0\}}\exp((x^{\prime}_{tj})^{\top}\mathbf{w}^{\star})}$
Then, by dividing the denominator and numerator by $\exp\left((x^{\prime}_{t0})^{\top}\mathbf{w}^{\star}\right)$, and defining $x_{ti}:=x^{\prime}_{ti}-x^{\prime}_{t0}$, we obtain the MNL probability in the form presented in (1) with $v_{0}=\exp(0)=1$. Note that this division does not change the probability.
B.2 Self-concordant-like Function
Definition B.1 (Self-concordant-like function, Tran-Dinh et al. 27).
@@ B.2 Self-concordant-like Function
$\phi^{\prime\prime\prime}(s)=\frac{\sum_{i<j}(a_{i}-a_{j})^{2}e^{a_{i}s+b_{i}}e^{a_{j}s+b_{j}}\left[\sum_{k=0}^{n}(a_{i}+a_{j}-2a_{k})e^{a_{k}s+b_{k}}\right]}{\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)^{3}}.$ (B.1)
@@
$|a_{i}+a_{j}-2a_{k}|\leq\sqrt{6}\sqrt{a_{i}^{2}+a_{j}^{2}+a_{k}^{2}}\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|.$ (B.2)
@@
$\left|\sum_{k=0}^{n}(a_{i}+a_{j}-2a_{k})e^{a_{k}s+b_{k}}\right|\leq\sum_{k=0}^{n}\left|a_{i}+a_{j}-2a_{k}\right|e^{a_{k}s+b_{k}}\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|\sum_{k=0}^{n}e^{a_{k}s+b_{k}}.$ (B.3)
Plugging in (B.3) into (B.1), we obtain
$\phi^{\prime\prime\prime}(s)\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|\phi^{\prime\prime}(s).$ (B.4)
Now, we are ready to prove the proposition. For any $t\in[T]$, let $n=|S_{t}|$ and $[c_{1},\dots c_{n}]=[x_{ti_{1}},\dots,x_{ti_{n}}]$. Define a function $f\in\mathcal{C}^{3}:\mathbb{R}^{d}\rightarrow\ \mathbb{R}$ as $f(\bm{\theta}):=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{T}\bm{\theta}}\right)$. Let $\bm{\delta}\in\mathbb{R}^{d}$ and let $f(\bm{\theta}+s\bm{\delta})=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{\top}\bm{\theta}+sc_{i}^{\top}\bm{\delta}}\right)=\log\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)=\phi(s)$, where $a_{i}=c_{i}^{\top}\bm{\delta}$, $b_{i}=c_{i}^{\top}\bm{\theta}$ for $i=1,\dots,n$, and $a_{i}=0$ and $b_{i}=\log v_{0}$ for $i=0$. Then, by (B.4), we get
@@
where the last inequality holds due to Assumption  1 that $\|c_{i}\|_{2}=\|x_{tj_{i}}\|_{2}\leq 1$. Then, by Definition B.1, $f$ is $3\sqrt{2}$-self-concordant-like. Since $\ell_{t}$ is the sum of $f$ and a linear operator, which has third derivatives equal to zero, it follows that $\ell_{t}$ is also $3\sqrt{2}$-self-concordant-like function. ∎
Remark B.1.
@@
Lemma B.1 (Theorem 3 of Tran-Dinh et al. 27).
@@
Appendix C Proof of Theorem 1
In this section, we provide the proof of Theorem 1. The proof structure is similar to the one presented in Chen et al. [9]. However, unlike their approach, we explicitly derive a bound that includes $v_{0}$. Furthermore, by establishing a tighter upper bound for the KL divergence (Lemma C.2), we derive a bound that is tighter than the one provided by Chen et al. [9].
C.1 Adversarial Construction and Bayes Risk
@@
$\geq\max_{\mathbf{w}_{V}\in\mathcal{W}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}R(S^{\star},\mathbf{w}_{V})-R(S_{t},\mathbf{w}_{V})$
@@
$=\frac{1}{|\mathcal{V}_{d/4}|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in S_{t}}p(i|S_{t},\mathbf{w}_{V})\right].$ (C.1)
@@
C.2 Main Proof of Theorem 1
@@
For any sequence of assortments $\{S_{t}\}_{t=1}^{T}$ produced by policy $\pi$, denote an alternative sequence $\{\tilde{S}_{t}\}_{t=1}^{T}$ that provably enjoys less regret under parameterization $\mathbf{w}_{V}$, for simplicity of presentation.
Let $x_{U_{1}},\dots,x_{U_{L}}$ be the distinct feature vectors contained in assortments $S_{t}$ (if $S_{t}=\emptyset$, then one may choose an arbitrary feature $x_{U}$) with $U_{1},\dots,U_{L}\in\mathcal{V}_{d/4}$. Let $U^{\star}$ be the subset among $U_{1},\dots,U_{L}$ that maximizes $x_{U}^{\top}\mathbf{w}_{V}$, i.e., $U^{\star}\in\mathop{\mathrm{argmax}}_{U\in\{U_{1},\dots,U_{L}\}}x_{U}^{\top}\mathbf{w}_{V}$, where $\mathbf{w}_{V}$ is the underlying parameter. Then, define $\tilde{S}_{t}$ as the assortment consisting of all $K$ items corresponding to feature $x_{U^{\star}}$, i.e., $\tilde{S}_{t}=\{\underbrace{x_{U^{\star}},\dots,x_{U^{\star}}}_{K}\}$.
@@
Proposition C.1 (Proposition 1 in Chen et al. 9).
@@
Hence, it is sufficient to bound $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$ instead of $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in S_{t}}p(i|S_{t},\mathbf{w}_{V})$.
@@
For any $j\in V$, define random variables $\tilde{M}_{j}:=\sum_{t=1}^{T}\mathbf{1}\{j\in\tilde{U}_{t}\}$. Then, by Lemma C.1, for all $V\in\mathcal{V}_{d/4}$, we have
$\mathbb{E}_{V}\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{4}-\sum_{j\in V}\mathbb{E}_{V}[\tilde{M}_{j}]\right).$ (C.2)
Furthermore, we define $\mathcal{V}_{d/4}^{(j)}:=\{V\in\mathcal{V}_{d/4}:j\in V\}$ and $\mathcal{V}_{d/4-1}:=\{V\subseteq[d]:|V|=d/4-1\}$. By taking the average of both sides of Equation (C.2) with respect to all $V\in\mathcal{V}_{d/4}$, we obtain
@@
$\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{6}-\max_{V\in\mathcal{V}_{d/4-1}}\sum_{j\notin V}\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|\right).$ (C.3)
Now we bound the term $\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|$ in (C.3) for any $V\in\mathcal{V}_{d/4-1}$. For simplicity, let $P=\mathbb{P}_{V}$ and $Q=\mathbb{P}_{V\cup\{j\}}$ denote the laws under $\mathbf{w}_{V}$ and $\mathbf{w}_{V\cup j}$, respectively. Then, we have
@@
$\leq T\cdot\|P-Q\|_{\operatorname{TV}}\leq T\cdot\sqrt{\frac{1}{2}\operatorname{KL}(P\|Q)},$ (C.4)
@@
Therefore, combining (C.3), (C.4), and Lemma C.2, we have
@@
C.3 Proofs of Lemmas for Theorem 1
C.3.1 Proof of Lemma C.1
Proof of Lemma C.1.
Let $x=x_{V}$ and $\hat{x}=x_{\tilde{U}_{t}}$ be the corresponding context vectros. Then, we have

### v6

@@ 8 Conclusion
Acknowledgements
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. 2022R1C1C1006859, 2022R1A4A1030579, and RS-2023-00222663) and by AI-Bio Research Grant through Seoul National University.
Appendix A Further Related Work
In this section, we discuss additional related works that complement Section 2. For simplicity, we consider only the dependence on the number of rounds $t$ for a computation cost in big-$\mathcal{O}$ notation.
[…]
Multinomial Logistic (MNL) Bandits. There are two main approaches to multinomial logistic (MNL) bandits: the multiple-parameter choice model and the single-parameter choice model. In the multiple-parameter choice model, the learner estimates parameters for each choice in the assortment ($\mathbf{w}_{1}^{\star},\dots,\mathbf{w}_{K}^{\star}$) with a shared context $x_{t}$. In this setting, Amani and Thrampoulidis [8] proposed a feasible algorithm that achieves a regret upper bound of $\tilde{\mathcal{O}}(dK\sqrt{\kappa T})$ with a $\mathcal{O}(t)$ computation cost. They also proposed an intractable algorithm that achieves an improved regret of $\tilde{\mathcal{O}}(dK^{3/2}\sqrt{T})$. Zhang and Sugiyama [52] introduced a computationally and statistically efficient algorithm that obtains a regret of $\tilde{\mathcal{O}}(dK\sqrt{T})$. Recently, Lee et al. [32] further improved the regret by a factor of $\sqrt{K}$, achieving $\tilde{\mathcal{O}}(d\sqrt{KT})$ regret. In the multiple-parameter case, the regret’s dependence on $K$ is unavoidable since the number of unknown parameters depends on $K$.
On the other hand, the single-parameter choice model, closely related to ours, shares the parameter $\mathbf{w}^{\star}$ cross the choices, with varying contexts for each choice. The learner offers a set of items $S_{t}$, with $|S_{t}|\leq K$ at each round. This setting involves a combinatorial optimization to choose the assortment $S_{t}$, making it more challenging to devise a tractable algorithm. As extensively discussed in Section 2, no previous studies have definitively confirmed whether the existing lower or upper bounds are tight. As shown in Table 1, many studies have presented their results in inconsistent settings with varying reward structures and values of $v_{0}$, adding to the ambiguity about the bounds’ optimality. In this paper, we address these issues by bridging the gap between the lower and upper bounds of regret through a careful categorization of the settings. We propose an algorithm that is both provably optimal, up to logarithmic factors, and computationally efficient, significantly enhancing the theoretical and practical understanding of MNL bandits.
Generalized Linear Bandits. In generalized linear bandits [22, 26, 34, 2, 30, 27, 28, 33], the expected rewards are modeled using a generalized linear model. These problems generalize logistic bandits by incorporating a general exponential family link function instead of the logistic link function. The algorithms proposed for generalized linear bandits also exhibit a dependence on the nonlinear term $\kappa$. However, our problem setting (single-parameter MNL bandits) considers a more complex state space where multiple arms are pulled simultaneously.
Combinatorial Bandits. Another related stream of literature is combinatorial bandits [12, 44, 29, 53, 45, 35], particularly top-$k$ combinatorial bandits [45]. In top-$k$ combinatorial bandits, the decision set includes all subsets of size $k$ out of $n$ arms, and the reward for each action is the sum of the rewards of the $k$ selected arms. In this framework, the rewards are assumed to be independent of the entire set of arms played in round $t$. In contrast, in our setting, the reward for each individual arm depends on the whole set of arms played.
Recently, Choi et al. [17], Cao and Sun [11] have considered the cascading assortment bandit problem, which encompasses the MNL bandit problem as a special case where the cascading length is $1$. However, these works do not strictly encompass our results. Choi et al. [17] only consider uniform rewards, and achieve a regret upper bound of $\tilde{\mathcal{O}}(d\sqrt{T})$, which avoids dependence on both the cascading length and $\kappa$. When the cascading length is $1$, our result (Theorem 2) improves upon theirs by a factor of $\sqrt{K}$. Moreover, their computation cost per round is $\mathcal{O}(t)$ since they employ MLE to estimate the parameter. Cao and Sun [11] consider non-uniform rewards, and achieve a regret upper bound of $\tilde{\mathcal{O}}(h^{2}d\sqrt{MT})$, where $M$ is the cascading length and $h\geq\frac{p(i|S,\mathbf{w})}{p(i|S,\mathbf{w^{\prime}})}$ for all $\mathbf{w},\mathbf{w}^{\prime}\in\mathcal{W}$, $S\in\mathcal{S}$, and $i\in S\cup\{0\}$. However, their regret bound still suffers from a harmful dependence on $h^{2}$, which can be exponentially large.
Appendix B Notation
We denote $T$ as the total number of rounds and $t\in[T]$ as the current round. We denote $N$ as the total number of items, $K$ as the maximum size of assortments, and $d$ as the dimension of feature vectors.
@@ Appendix B Notation
We offer a Table B.1 for convenient reference.
Table B.1: Symbols
@@
$\beta_{t}(\delta)$ $:=\mathcal{O}\left(\sqrt{d}\log t\log K\right)$, confidence radius at round $t$
$\tilde{R}_{t}(S)$ $:=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})}$, optimistic expected revenue for the assortment $S$ at round $t$
Appendix C Properties of MNL function
@@
C.1 Attraction parameter for Outside Option: $v_{0}=\Theta(1)$ is Common in Contextual MNL Bandits
In this subsection, we explain why the assumption that $v_{0}=\Theta(1)$ is made without loss of generality. Let the original feature vectors be $x^{\prime}_{ti}\in\mathbb{R}^{d}$ for every item $i\in[N]$. Suppose that a context for the outside option $x_{t0}^{\prime}$ is given and the probability of choosing any item $i\in S_{t}\cup\{0\}$ is defined as
$p_{t}(i|S_{t},\mathbf{w}^{\star})=\frac{\exp((x^{\prime}_{ti})^{\top}\mathbf{w}^{\star})}{\sum_{j\in S_{t}\cup\{0\}}\exp((x^{\prime}_{tj})^{\top}\mathbf{w}^{\star})}.$
Then, by dividing the denominator and numerator by $\exp\left((x^{\prime}_{t0})^{\top}\mathbf{w}^{\star}\right)$, and defining $x_{ti}:=x^{\prime}_{ti}-x^{\prime}_{t0}$, we obtain the MNL probability in the form presented in (1) with $v_{0}=\exp(0)=1$. Note that this division does not change the probability. Therefore, $v_{0}=\Theta(1)$ is natural and common in contextual MNL bandit literature.
C.2 Self-concordant-like Function
Definition C.1 (Self-concordant-like function, Tran-Dinh et al. 48).
@@ C.2 Self-concordant-like Function
$\phi^{\prime\prime\prime}(s)=\frac{\sum_{i<j}(a_{i}-a_{j})^{2}e^{a_{i}s+b_{i}}e^{a_{j}s+b_{j}}\left[\sum_{k=0}^{n}(a_{i}+a_{j}-2a_{k})e^{a_{k}s+b_{k}}\right]}{\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)^{3}}.$ (C.1)
@@
$|a_{i}+a_{j}-2a_{k}|\leq\sqrt{6}\sqrt{a_{i}^{2}+a_{j}^{2}+a_{k}^{2}}\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|.$ (C.2)
@@
$\left|\sum_{k=0}^{n}(a_{i}+a_{j}-2a_{k})e^{a_{k}s+b_{k}}\right|\leq\sum_{k=0}^{n}\left|a_{i}+a_{j}-2a_{k}\right|e^{a_{k}s+b_{k}}\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|\sum_{k=0}^{n}e^{a_{k}s+b_{k}}.$ (C.3)
Plugging in (C.3) into (C.1), we obtain
$\phi^{\prime\prime\prime}(s)\leq 3\sqrt{2}\max_{i=0,\dots,n}|a_{i}|\phi^{\prime\prime}(s).$ (C.4)
Now, we are ready to prove the proposition. For any $t\in[T]$, let $n=|S_{t}|$ and $c_{1}=x_{ti_{1}},c_{2}=x_{ti_{2}},\dots,c_{n}=x_{ti_{n}}$. Define a function $f\in\mathcal{C}^{3}:\mathbb{R}^{d}\rightarrow\ \mathbb{R}$ as $f(\bm{\theta}):=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{\top}\bm{\theta}}\right)$. Let $\bm{\delta}\in\mathbb{R}^{d}$ and let $f(\bm{\theta}+s\bm{\delta})=\log\left(v_{0}+\sum_{i=1}^{n}e^{c_{i}^{\top}\bm{\theta}+sc_{i}^{\top}\bm{\delta}}\right)=\log\left(\sum_{i=0}^{n}e^{a_{i}s+b_{i}}\right)=\phi(s)$, where $a_{i}=c_{i}^{\top}\bm{\delta}$, $b_{i}=c_{i}^{\top}\bm{\theta}$ for $i=1,\dots,n$, and $a_{i}=0$ and $b_{i}=\log v_{0}$ for $i=0$. Then, by (C.4), we get
@@
where the last inequality holds due to Assumption  1 that $\|c_{i}\|_{2}=\|x_{tj_{i}}\|_{2}\leq 1$. Then, by Definition C.1, $f$ is $3\sqrt{2}$-self-concordant-like. Since $\ell_{t}$ is the sum of $f$ and a linear operator, which has third derivatives equal to zero, it follows that $\ell_{t}$ is also $3\sqrt{2}$-self-concordant-like function. ∎
Remark C.1.
@@
Lemma C.1 (Theorem 3 of Tran-Dinh et al. 48).
@@
Appendix D Proof of Theorem 1
In this section, we provide the proof of Theorem 1. The proof structure is similar to the one presented in Chen et al. [14]. However, unlike their approach, we explicitly derive a bound that includes $v_{0}$. Furthermore, by establishing a tighter upper bound for the KL divergence (Lemma D.2), we derive a bound that is tighter than the one provided by Chen et al. [14].
D.1 Adversarial Construction and Bayes Risk
@@
$\geq\max_{\mathbf{w}_{V}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}R(S^{\star},\mathbf{w}_{V})-R(S_{t},\mathbf{w}_{V})$
@@
$=\frac{1}{|\mathcal{V}_{d/4}|}\sum_{V\in\mathcal{V}_{d/4}}\mathbb{E}_{\mathbf{w}_{V}}^{\pi}\sum_{t=1}^{T}\left[\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in S_{t}}p(i|S_{t},\mathbf{w}_{V})\right].$ (D.1)
@@
D.2 Main Proof of Theorem 1
@@
For any sequence of assortments $\{S_{t}\}_{t=1}^{T}$ produced by policy $\pi$, we denote an alternative sequence $\{\tilde{S}_{t}\}_{t=1}^{T}$ that provably enjoys less regret under parameterization $\mathbf{w}_{V}$.
Let $x_{U_{1}},\dots,x_{U_{L}}$ be the distinct feature vectors contained in assortments $S_{t}$ (if $S_{t}=\emptyset$, then one may choose an arbitrary feature $x_{U}$) with $U_{1},\dots,U_{L}\in\mathcal{V}_{d/4}$. Let $U^{\star}$ be the subset among $U_{1},\dots,U_{L}$ that maximizes $x_{U}^{\top}\mathbf{w}_{V}$, i.e., $U^{\star}\in\mathop{\mathrm{argmax}}_{U\in\{U_{1},\dots,U_{L}\}}x_{U}^{\top}\mathbf{w}_{V}$, where $\mathbf{w}_{V}$ is the underlying parameter. Then, we define $\tilde{S}_{t}$ as the assortment consisting of all $K$ items corresponding to feature $x_{U^{\star}}$, i.e., $\tilde{S}_{t}=\{\underbrace{x_{U^{\star}},\dots,x_{U^{\star}}}_{K}\}$.
@@
Proposition D.1 (Proposition 1 in Chen et al. 14).
@@
Proposition D.1 implies that $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in S_{t}}p(i|S_{t},\mathbf{w}_{V})\geq\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$. Hence, it is sufficient to bound $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})$ instead of $\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in S_{t}}p(i|S_{t},\mathbf{w}_{V})$.
@@
For any $j\in V$, define random variables $\tilde{M}_{j}:=\sum_{t=1}^{T}\mathbf{1}\{j\in\tilde{U}_{t}\}$. Then, by Lemma D.1, for all $V\in\mathcal{V}_{d/4}$, we have
$\mathbb{E}_{V}\sum_{i\in S^{\star}}p(i|S^{\star},\mathbf{w}_{V})-\sum_{i\in\tilde{S}_{t}}p(i|\tilde{S}_{t},\mathbf{w}_{V})\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{4}-\sum_{j\in V}\mathbb{E}_{V}[\tilde{M}_{j}]\right).$ (D.2)
Furthermore, we define $\mathcal{V}_{d/4}^{(j)}:=\{V\in\mathcal{V}_{d/4}:j\in V\}$ and $\mathcal{V}_{d/4-1}:=\{V\subseteq[d]:|V|=d/4-1\}$. By taking the average of both sides of Equation (D.2) with respect to all $V\in\mathcal{V}_{d/4}$, we obtain
@@
$\geq\frac{v_{0}K}{(v_{0}+Ke)^{2}}\cdot\frac{\epsilon}{2\sqrt{d}}\left(\frac{dT}{6}-\max_{V\in\mathcal{V}_{d/4-1}}\sum_{j\notin V}\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|\right).$ (D.3)
Now we bound the term $\left|\mathbb{E}_{V\cup\{j\}}[\tilde{M}_{j}]-\mathbb{E}_{V}[\tilde{M}_{j}]\right|$ in (D.3) for any $V\in\mathcal{V}_{d/4-1}$. For simplicity, let $P=\mathbb{P}_{V}$ and $Q=\mathbb{P}_{V\cup\{j\}}$ denote the laws under $\mathbf{w}_{V}$ and $\mathbf{w}_{V\cup j}$, respectively. Then, we have
@@
$\leq T\cdot\|P-Q\|_{\operatorname{TV}}\leq T\cdot\sqrt{\frac{1}{2}\operatorname{KL}(P\|Q)},$ (D.4)
@@
Therefore, combining (D.3), (D.4), and Lemma D.2, we have
@@
D.3 Proofs of Lemmas for Theorem 1
D.3.1 Proof of Lemma D.1
Proof of Lemma D.1.
Let $x=x_{V}$ and $\hat{x}=x_{\tilde{U}_{t}}$ be the corresponding context vectors. Then, we have